OpenStudy (anonymous):

x^2=10x-20 (x-5)^2=2

OpenStudy (saifoo.khan):

factor? or solutions?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[\pm\]

OpenStudy (saifoo.khan):

(x-5)^2-5 = 0

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

\[(x-5)^2=2\] \[x-5=\pm\sqrt{2}\] \[x=5\pm\sqrt{2}\]

OpenStudy (anonymous):

first one is \[x^2-10x=-20\] \[(x-5)^2=-20+25\] \[(x-5)^2=5\] \[x-5=\pm\sqrt{5}\] \[x=5\pm\sqrt{5}\]